40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 7 23:24:41 2014, Jenne, Update, General, IFO plan, list o' things to do 
    Reply  Wed Jan 8 13:46:39 2014, Jenne, Update, General, IFO plan, PRM baffle PRMcage_6Sept2012.JPG
    Reply  Wed Jan 8 17:53:26 2014, manasa, Update, General, IFO plan, IPANG telescope IPANG_layout.pngIPANG.png
       Reply  Thu Jan 9 17:21:45 2014, rana, Update, General, IFO plan, IPANG telescope 
    Reply  Thu Jan 9 10:34:58 2014, Steve, Update, PSL, PSL pointing changes in pitch Lfnfdnc.png24hPSLpointing.png
       Reply  Fri Jan 10 10:28:03 2014, Steve, Update, PSL, PSL pointing changes  clamped.png
Message ID: 9545     Entry time: Fri Jan 10 10:28:03 2014     In reply to: 9542
Author: Steve 
Type: Update 
Category: PSL 
Subject: PSL pointing changes  

 

I looked at IOO QPDs again. QPD_POS was clamped by one screw. Dog clamp was added on the unclamped side.

QPD_ANG chassis has no isolation to optical table..._POS has.  QPD_ANG  base was tightened also.

Both QPDs moved a little bit but I did not centered them.  The spot sizes are 2-3 mm  They should be smaller.

How ever, we still can not explane the pitch movement of the IOO beam

 

Razor beam dumps were labeled at the AP table.

 

The 40m roof was cleaned from leafs this morning.

 

 

Attachment 1: clamped.png  30 kB  | Show | Show all
ELOG V3.1.3-